OpenStudy (anonymous):

What are the solutions of the equation? c^2-10c=0

OpenStudy (anonymous):

0 and 10

OpenStudy (anonymous):

c^2-10c=0 c(c-10)=0 Either c has to be 0 or c-10 has to be 0. c=0 or c-10=0 c=0 or c=10

OpenStudy (anonymous):

hey romero, what happened to showing ur work??

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

no different than nightwill